# Fairness-Aware Domain Generalization under Covariate and Dependence Shifts

### v1

@@ 2 Preliminaries
Specifically, when $p_{1}=\mathbb{P}(Z=1)$ and $p_{1}=\mathbb{P}(Z=1,Y=1)$, the fairness notion $\rho(\hat{Y},Z)$ is defined as the difference of demographic parity (DP) and the difference of equalized odds (EO), respectively [15]. In the paper, we will present the results under DP, while the framework can be generalized to multi-class, multi-sensitive attributes and other fairness notions. Strictly speaking, a classifier $f$ is fair over subgroups if it satisfies $\rho(\hat{Y},Z)=0$.
Problem Setting. We consider a set of data domains $\mathcal{E}$, where each domain $e\in\mathcal{E}$ corresponds to a distinct data subset $\mathcal{D}^{e}=\{(\mathbf{x}^{e}_{i},z^{e}_{i},y^{e}_{i})\}_{i=1}^{|\mathcal{D}^{e}|}$ sampled i.i.d. from $\mathbb{P}(X^{e},Z^{e},Y^{e})$. Given a dataset $\mathcal{D}=\{\mathcal{D}^{e}\}_{e\in\mathcal{E}}$, it is partitioned into multiple source domains $\mathcal{E}_{s}\subset\mathcal{E}$ and unknown target domains which are inaccessible during training. Therefore, given samples from finite source domains $\mathcal{E}_{s}$, the goal of fairness-aware domain generalization problems is to learn a classifier $f\in\mathcal{F}$ that is generalizable across all possible domains.
Problem 1 (Fairness-aware Domain Generalization).
Let $\mathcal{E}_{s}\subset\mathcal{E}$ be a finite subset of source domains and assume that, for each $e\in\mathcal{E}_{s}$, we have access to its corresponding dataset $\mathcal{D}^{e}=\{(\mathbf{x}^{e}_{i},z^{e}_{i},y^{e}_{i})\}_{i=1}^{|\mathcal{D}^{e}|}$ sampled i.i.d from $\mathbb{P}(X^{e},Z^{e},Y^{e})$. Given a classifier set $\mathcal{F}$ and a loss function $\ell:\mathcal{Y\times Y}\rightarrow\mathbb{R}$, the goal is to learn a fair classifier $f\in\mathcal{F}$ for any $\mathcal{D}^{e}\in\mathcal{D}_{s}=\{\mathcal{D}^{e}\}_{e\in\mathcal{E}_{s}}$ that minimizes the worst-case risk over all domains in $\mathcal{E}$ satisfying a group fairness constraint:
$\min_{f\in\mathcal{F}}\>\max_{e\in\mathcal{E}}\>\mathbb{E}_{\mathbb{P}(X^{e},Z^{e},Y^{e})}\ell(f(X^{e}),Y^{e}),\quad\text{subject to}\>\>\rho(f(X^{e}),Z^{e})=0.$ (2)
[…]
Another challenge is how closely the data distributions in unknown target domains match those in the observed source domains. In [7], there are two forms of distribution shifts: concept shift, where the instance conditional distribution $\mathbb{P}(Y^{e}|X^{e},Z^{e})$ varies across different domains, and covariate shift, where the marginal distributions over instance $\mathbb{P}(X^{e})$ are various. Yet, neither of these shifts captures the degree to which the distribution shifts with regard to model fairness. Therefore, we introduce a novel variation, dependence shift, where the dependence $\rho(Y^{e},Z^{e})$ between sensitive attributes and sample classes differs across domains.
Definition 2 (Covariate Shift [12] and Dependence Shift).
In Prob. 1, covariate shift occurs when environmental variation is attributed to disparities in the marginal distributions across instances $\{\mathbb{P}(X^{e})\}_{e\in\mathcal{E}}$. On the other hand, Prob. 1 exhibits a dependence shift when environmental variation arises from alterations in the sensitive dependence $\rho(Y^{e},Z^{e})$.
3 Fairness-aware Disentangled Domain Generalization
DG tasks can generally be characterized by one form of the three distribution shifts. In this paper, we restrict the scope of our framework to focus on Prob. 1 in which inter-domain variation is due to covariate shift and dependence shift simultaneously through an underlying transformation model $T$.
Assumption 1 (Transformation Model).
[…]
Under Assump. 1, a data subset $\mathcal{D}^{e_{j}}\in\mathcal{D}$ of domain $e_{j}$ can be regarded as generated from another data subset $\mathcal{D}^{e_{i}}$ through the transformation model $T$ by altering $(X^{e_{i}},Z^{e_{i}})$ to $(X^{e_{j}},Z^{e_{j}})$. Building upon the insights from existing DG literature [13, 16, 17], we define $T$ with a specific emphasis on disentangling the variation in data features across domains into latent spaces with three factors. For specific information about the design of $T$ and the learning algorithm, please refer to Sec. 4.
Assumption 2 (Multiple Latent Factors).
Given dataset $\mathcal{D}^{e}=\{(\mathbf{x}^{e}_{i},z^{e}_{i},y^{e}_{i})\}_{i=1}^{|\mathcal{D}^{e}|}$ sampled i.i.d. from $\mathbb{P}(X^{e},Z^{e},Y^{e})$ in domain $e\in\mathcal{E}$, we assume that each instance $\mathbf{x}^{e}_{i}$ is generated from
  • a latent content factor $\mathbf{c}\in\mathcal{C}$, where $\mathcal{C}=\{\mathbf{c}_{y=0},\mathbf{c}_{y=1}\}$ refers to a content space;
  • a latent sensitive factor $\mathbf{a}\in\mathcal{A}$, where $\mathcal{A}=\{\mathbf{a}_{z=1},\mathbf{a}_{z=-1}\}$ refers to a sensitive space;
  • a latent style factor $\mathbf{s}^{e}$, where $\mathbf{s}^{e}$ is specific to the individual domain $e$.
We assume that the content and sensitive factors in $\mathcal{C}$ and $\mathcal{A}$ do not change across domains. Each domain $e$ over $\mathbb{P}(X^{e},Z^{e},Y^{e})$ is represented by a style factor $\mathbf{s}^{e}$ and the dependence score $\rho^{e}=\rho(Y^{e},Z^{e})$, denoted $e:=(\mathbf{s}^{e},\rho^{e})$, where $\mathbf{s}^{e}$ and $\rho^{e}$ are unique to the domain $e$.
Note that Assump. 2 is similarly related to the one made in [13, 12, 18, 19]. In our paper, with a focus on group fairness, we expand upon the assumptions of existing works by introducing three latent factors. Because we assume the instance conditional distribution $\mathbb{P}(Y|X,Z)$ remains consistent across domains (i.e., there is no concept shift taking place), under Assumps. 1 and 2, if two instances $(\mathbf{x}^{e_{i}},z^{e_{i}},y)$ and $(\mathbf{x}^{e_{j}},z^{e_{j}},y)$ where $e_{i},e_{j}\in\mathcal{E},i\neq j$ share the same class label, then the latter instance can be reconstructed from the former using $T$. Specifically, $T$ constructs $\mathbf{x}^{e_{j}}$ using the content factor of $\mathbf{x}^{e_{i}}$, the sensitive factor of $\mathbf{x}^{e_{j}}$, and the style factor of $\mathbf{x}^{e_{j}}$. Additionally, $T$ constructs $z^{e_{j}}$ by employing the sensitive factor of $\mathbf{x}^{e_{j}}$. For fairness-aware invariant learning, we make the following assumption.
@@ 3 Fairness-aware Disentangled Domain Generalization
We assume that inter-domain variation is characterized by the covariate shift and dependence shift in $\mathbb{P}(X^{e})$ and $\rho(Y^{e},Z^{e}),\forall e\in\mathcal{E}$. As a consequence, we assume that $\mathbb{P}(Y^{e}|X^{e},Z^{e})$ is stable across domains. Given a domain transformation function $T$, for any $\mathbf{x}\in\mathcal{X}$, $z\in\mathcal{Z}$, and $y\in\mathcal{Y}$, it holds that
$\mathbb{P}(Y^{e_{i}}=y|X^{e_{i}}=\mathbf{x}^{e_{i}},Z^{e_{i}}=z^{e_{i}})=\mathbb{P}(Y^{e_{j}}=y|(X^{e_{j}},Z^{e_{j}})=T(\mathbf{x}^{e_{i}},z^{e_{i}},e_{j})),\>\forall e_{i},e_{j}\in\mathcal{E},i\neq j$
In Assump. 3, the domain shift captured by $T$ would characterize the mapping from the underlying distributions $\mathbb{P}(X^{e_{i}})$ and $\rho(Y^{e_{i}},Z^{e_{i}})$ over $\mathcal{D}^{e_{i}}$ to the distribution $\mathbb{P}(X^{e_{j}})$ and $\rho(Y^{e_{j}},Z^{e_{j}})$ of samples from a different data domain $\mathcal{D}^{e_{j}}$, respectively. With this in mind and under Assump. 3, we introduce a new definition of fairness-aware invariance with respect to the variation captured by $T$ and satisfying the group fair constraint introduced in Defn. 1.
@@
Given a transformation model $T$, a fairness-aware classifier $f\in\mathcal{F}$ is domain invariant if it holds
[…]
almost surely when $(\mathbf{x}^{e_{j}},z^{e_{j}})=T(\mathbf{x}^{e_{i}},z^{e_{i}},e_{j})$, $\mathbf{x}^{e_{i}}\sim\mathbb{P}(X^{e_{i}})$, $\mathbf{x}^{e_{j}}\sim\mathbb{P}(X^{e_{j}})$, and $e_{i},e_{j}\in\mathcal{E}$.
Defn. 3 is crafted to enforce invariance on the predictions generated by $f$ directly. We expect a prediction to yield the same prediction for any realization of data under $T$ while being aware of group fairness.
Problem 2 (Fairness-aware Disentanglement for Domain Generalization).
Under Defn. 3 and Assump. 3, if we restrict $\mathcal{F}$ of Prob. 1 to the set of invariant fairness-aware classifiers, the Prob. 1 is equivalent to the following problem
$P^{\star}\triangleq\min_{f\in\mathcal{F}}\>\>$ $R(f)\triangleq\mathbb{E}_{\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}})}\ell(f(X^{e_{i}}),Y^{e_{i}})$ (4)
subject to $f(\mathbf{x}^{e_{i}})=f(\mathbf{x}^{e_{j}}),\quad\rho(f(X^{e_{i}}),Z^{e_{i}})=\rho(f(X^{e_{j}}),Z^{e_{j}})=0$
where $\mathbf{x}^{e_{i}}\sim\mathbb{P}(X^{e_{i}})$, $\mathbf{z}^{e_{i}}\sim\mathbb{P}(Z^{e_{i}})$, $(\mathbf{x}^{e_{j}},z^{e_{j}})=T(\mathbf{x}^{e_{i}},z^{e_{i}},e_{j})$, $\forall e_{i},e_{j}\in\mathcal{E}$, $i\neq j$.
Similar to [12], Prob. 2 is not a composite optimization problem. Moreover, acquiring domain labels is often expensive or even unattainable, primarily due to privacy concerns. Consequently, under the assumptions of disentanglement-based invariance and domain shift, Problem 1 can be approximated to Problem 2 by removing the max operator. Furthermore, Prob. 2 offers a new and theoretically-principled perspective on Prob. 1, when data varies from domain to domain with respect to an underlying transformation model $T$. To optimize Prob. 2 is challenging because
  • The strict equality constraints in Prob. 2 are difficult to enforce in practice;
  • Enforcing constraints on deep networks is known to be a challenging problem due to non-convexity. Simply transforming them to regularization cannot guarantee satisfaction for constrained problems;
  • As we have incomplete access to all domains $\mathcal{E}$ or $\mathbb{P}(X,Z,Y)$, it limits the ability to enforce fairness-aware $T$-invariance and further makes it hard to estimate $R(f)$.
@@
$P^{\star}(\gamma_{1},\gamma_{2})\triangleq\min_{f\in\mathcal{F}}R(f)\quad\text{subject to}\quad\delta^{e_{i},e_{j}}(f)\leq\gamma_{1},\epsilon^{e_{i}}(f)\leq\frac{\gamma_{2}}{2}\text{ and }\epsilon^{e_{j}}(f)\leq\frac{\gamma_{2}}{2}$ (5)
@@
$\delta^{e_{i},e_{j}}(f)\triangleq\mathbb{E}_{\mathbb{P}(X,Z)}d\big{[}f(X^{e_{i}}),f(X^{e_{j}}=T(X^{e_{i}},Z^{e_{i}},e_{j}))\big{]},$ (6)
$\epsilon^{e_{i}}(f)\triangleq\rho(f(X^{e_{i}}),Z^{e_{i}}),\quad\epsilon^{e_{j}}(f)\triangleq\rho(f(X^{e_{j}}),Z^{e_{j}})$ (7)
and $\forall e_{i},e_{j}\in\mathcal{E}$, $i\neq j$. Here, $\gamma_{1},\gamma_{2}>0$ are constants controlling the extent of relaxation and $d[\cdot]$ is a distance metric, e.g., KL-divergence. When $\gamma_{1}=\gamma_{2}=0$, Eqs. 4 and 5 are equivalent.
Theorem 1 (Fairness Upper Bound of the Unseen Target Domain).
In accordance with Defns. 1 and 7, for any domain $e\in\mathcal{E}$, the fairness dependence under instance distribution $\mathbb{P}(X^{e},Z^{e},Y^{e})$ with respect to the classifier $f\in\mathcal{F}$ is defined as:
$\epsilon^{e}(f)=\big{|}\mathbb{E}_{\mathbb{P}(X^{e},Z^{e})}g(f(X^{e}),Z^{e})\big{|}$
With observed source domains $\mathcal{E}_{s}$, the dependence at any unseen target domain $e_{t}\in\mathcal{E}\backslash\mathcal{E}_{s}$ is upper bounded. $D[\cdot]$ is the Jensen-Shannon distance [20] metric.
$\epsilon^{e_{t}}(f)\leq$ $\frac{1}{|\mathcal{E}_{s}|}\sum_{e_{i}\in\mathcal{E}_{s}}\epsilon^{e_{i}}(f)+\sqrt{2}\min_{e_{i}\in\mathcal{E}_{s}}D\big{[}\mathbb{P}(X^{e_{t}},Z^{e_{t}},Y^{e_{t}}),\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}})\big{]}$
$+\sqrt{2}\max_{e_{i},e_{j}\in\mathcal{E}_{s}}D\big{[}\mathbb{P}(X^{e_{i}},Z^{e_{i}},Y^{e_{i}}),\mathbb{P}(X^{e_{j}},Z^{e_{j}},Y^{e_{j}}\big{]}$
where $D[\mathbb{P}_{1},\mathbb{P}_{2}]=\sqrt{\frac{1}{2}KL(\mathbb{P}_{1}||\frac{\mathbb{P}_{1}+\mathbb{P}_{2}}{2})+\frac{1}{2}KL(\mathbb{P}_{2}||\frac{\mathbb{P}_{1}+\mathbb{P}_{2}}{2})}$ is JS divergence defined based on KL divergence.
@@
$D^{\star}_{\xi,N,\mathcal{E}_{s}}(\gamma_{1},\gamma_{2})$ $\triangleq\max_{\lambda_{1}({e_{i},e_{j}}),\lambda_{2}({e_{i},e_{j}})}\min_{\bm{\theta}\in\Theta}\>\hat{R}(\bm{\theta})$ (8)
$+\frac{1}{|\mathcal{E}_{s}|}\sum_{{e_{i},e_{j}}\in\mathcal{E}_{s}}\Big{[}\lambda_{1}{(e_{i},e_{j})}\big{(}\hat{\delta}^{e_{i},e_{j}}(\bm{\theta})-\gamma_{1}\big{)}+\lambda_{2}{(e_{i},e_{j})}\big{(}\hat{\epsilon}^{e_{i}}(\bm{\theta})+\hat{\epsilon}^{e_{j}}(\bm{\theta})-\gamma_{2}\big{)}\Big{]}$
@@
where $\xi=\mathbb{E}_{\mathbb{P}(X)}||f(\mathbf{x})-\hat{f}(\mathbf{x},\bm{\theta})||_{\infty}>0$ is a constant bounding the difference between $f$ and its parameterized counterpart $\hat{f}:\mathcal{X}\times\Theta\rightarrow\mathbb{R}$ defined in the Definition 5.1 of [12]. $\lambda_{1}{(e_{i},e_{j})},\lambda_{2}{(e_{i},e_{j})}>0$ are dual variables. $\hat{R}(\bm{\theta})$, $\hat{\delta}^{e_{i},e_{j}}(\bm{\theta})$, $\hat{\epsilon}^{e_{i}}(\bm{\theta})$ and $\hat{\epsilon}^{e_{j}}(\bm{\theta})$ are the empirical counterparts of $R(\hat{f}(\cdot,\bm{\theta}))$, $\delta^{e_{i},e_{j}}(\hat{f}(\cdot,\bm{\theta}))$, $\epsilon^{e_{i}}(\hat{f}(\cdot,\bm{\theta}))$ and $\epsilon^{e_{j}}(\hat{f}(\cdot,\bm{\theta}))$, respectively. With such approximation on the dual problem in Eq. 8, the duality gap between $P^{\star}$ and $D^{\star}_{\xi,N,\mathcal{E}_{s}}(\gamma_{1},\gamma_{2})$ can be explicitly bounded.
Theorem 2 (Fairness-aware Data-dependent Duality Gap).
Given $\xi>0$, assuming $\{\hat{f}(\cdot,\bm{\theta}):\bm{\theta}\in\Theta\}\subseteq\mathcal{F}$ has finite VC-dimension, with $M$ datapoints sampled from $\mathbb{P}(X,Z,Y)$ we have
$|P^{\star}-D^{\star}_{\xi,N,\mathcal{E}_{s}}(\bm{\gamma})|\leq L||\bm{\gamma}||_{1}+\xi k(1+||\bm{\lambda}^{\star}_{p}||_{1})+O(\sqrt{\log(M)/M})$
where $\bm{\gamma}=[\gamma_{1},\gamma_{2}]^{T}$; $L$ is the Lipschitz constant of $P^{\star}(\gamma_{1},\gamma_{2})$; $k$ is a small universal constant defined in Proposition 3 of Appendix E; and $\bm{\lambda}^{\star}_{p}$ is the optimal dual variable for a perturbed version of Eq. 5.
The duality gap that arises when solving the empirical problem presented in Eq. 8 is minimal when the fairness-aware $T$-invariance in Defn. 3 margin $\bm{\gamma}$ is narrow, and the parametric space closely approximates $\mathcal{F}$. Proofs of Theorems 1 and 2 are provided in Appendix E.
4 An Effective Algorithm
Figure 2: An overview of data generation in synthetic domains via an underlying transformation model $T$. Data are augmented through $T$ based on invariant content factors and randomly sampled sensitive and style factors that encode synthetic domains. We demonstrate the concept using the ccMnist introduced in Sec. 5, where the domains are associated with various digit colors and the sensitive labels are determined by the background colors of each image.
Fairness-aware DG via $T$. Motivated by the theoretical insights in Sec. 3, we propose a simple but effective algorithm, namely FDDG. This algorithm consists of two stages. In the first stage, we train the transformation model $T$ using data from the source domains. In the second stage, we harness the power of $T$ to address the unconstrained dual optimization problem outlined in Eq. 8 through a series of primal-dual iterations.
Regarding the architecture of $T$, we expand upon the networks used in [18] by incorporating an additional output of the encoder $E:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{C}\times\mathcal{A}\times\mathcal{S}$ for sensitive factors $\mathbf{a}\in\mathcal{A}$ and including a sensitive classifier $h:\mathcal{A}\rightarrow\mathcal{Z}$. A generator (decoder) $G:\mathcal{C}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathcal{X}\times\mathcal{Z}$ is used to reconstruct instances from encoded latent factors. Following [18, 12], the transformation model $T$ is trained using data in the observed source domains by applying reconstruction of them. Detailed training process of $T$ is provided in Appendix C.
Given the finite number of observed source domains, to enhance the generalization performance for unseen target domains, the invariant classifier $\hat{f}$ is trained by expanding the dataset with synthetic domains. These synthetic domains are created by introducing random instance styles and random fair dependencies within the domain. As described in Fig. 2, the sensitive factor $\mathbf{a}^{e^{\prime}}$ and the style factor $\mathbf{s}^{e^{\prime}}$ are randomly sampled from their prior distributions $\mathcal{N}(0,\mathbf{I}_{a})$ and $\mathcal{N}(0,\mathbf{I}_{s})$, respectively. Along with the unchanged content factor $\mathbf{c}$, they are further passed through $G$ to generate a new instance within a novel domain. Under Assumps. 3 and 3, according to Eqs. 6 and 7, data augmented in synthetic domains are required to maintain invariance in terms of accuracy and fairness with the data in the corresponding original domains.
Walking Through Algorithm 1. Our proposed implementation is shown in Algorithm 1 to solve the empirical dual Eq. 8. In lines 15-20, we describe the DataAug procedure that takes an example $(\mathbf{x},z,y)$ as INPUT and returns an augmented example $(\mathbf{x}^{\prime},z^{\prime},y)$ from a new synthetic domain as OUTPUT. The augmented example has the same content factor as the input example but has different sensitive and style factors sampled from their associated prior distributions that encode a new synthetic domain. Lines 1-14 show the main training loop for FDDG. In line 6, for each example in the minibatch $\mathcal{B}$, we apply the procedure DataAug to generate an augmented example from a new synthetic domain described above. In line 7, we consider KL-divergence as the distance metric for $d[\cdot]$. All the augmented examples are stored in the set $\mathcal{B}^{\prime}$. The Lagrangian dual loss function is defined based on $\mathcal{B}$ and $\mathcal{B}^{\prime}$ in line 10. The primal parameters $\bm{\theta}$ and the dual parameters $\lambda_{1}$ and $\lambda_{2}$ are updated in lines 11-12.
Algorithm 1 Fair Disentangled Domain Generalization.
Require: Pretained encoder $E$, decoder $G$ and sensitive classifier $h$ within $T$.
@@ 4 An Effective Algorithm
3:         $\mathcal{L}_{cls}(\bm{\theta})=\frac{1}{m}\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}_{i},\bm{\theta}))$
4:         Initialize $\mathcal{L}_{inv}(\bm{\theta})=0$ and $\mathcal{B}^{\prime}=[\>]$
@@
Generate $(\mathbf{x}_{j},z_{j},y_{j})=\textsc{DataAug}(\mathbf{x}_{i},z_{i},y_{i})$ and add it to $\mathcal{B}^{\prime}$
@@
$\mathcal{L}_{inv}(\bm{\theta})+=\frac{1}{m}d[\hat{f}(\mathbf{x}_{i},\bm{\theta}),\hat{f}(\mathbf{x}_{j},\bm{\theta})]$

### v2

@@ 4.1 Learning the Transformation Model
$\mathcal{L}_{recon}^{data}=\mathbb{E}_{\mathbf{x}^{s}\sim\mathbb{P}^{s}_{X}}[\lVert G(E_{c}(\mathbf{x}^{s}),E_{a}(\mathbf{x}^{s}),E_{s}(\mathbf{x}^{s}))-\mathbf{x}^{s}\rVert_{1}]$
Factor reconstruction loss. Given latent factors $\mathbf{c}$, $\mathbf{a}$, and $\mathbf{s}^{s}$ encoded from a sample $\mathbf{x}^{s}$, they are encouraged to be reconstructed through some latent factors randomly sampled from the prior Gaussian distributions.
$\mathcal{L}_{recon}^{factor}=\>$ $\mathbb{E}_{\mathbf{c}\sim\mathbb{P}_{C},\mathbf{a}\sim\mathbb{P}_{A},\mathbf{s}^{s}\sim\mathbb{P}_{S}}[\lVert E_{c}(G(\mathbf{c,a},\mathbf{s}^{s}))-\mathbf{c}\rVert_{1}]$
$+\>\mathbb{E}_{\mathbf{c}\sim\mathbb{P}_{C},\mathbf{a}\sim\mathcal{N}(0,\mathbf{I}_{a}),\mathbf{s}^{s}\sim\mathbb{P}_{S}}[\lVert E_{a}(G(\mathbf{c,a},\mathbf{s}^{s}))-\mathbf{a}\rVert_{1}]$
$+\>\mathbb{E}_{\mathbf{c}\sim\mathbb{P}_{C},\mathbf{a}\sim\mathbb{P}_{A},\mathbf{s}^{s}\sim\mathcal{N}(0,\mathbf{I}_{s})}[\lVert E_{s}(G(\mathbf{c,a},\mathbf{s}^{s}))-\mathbf{s}\rVert_{1}]$
where $\mathbb{P}_{C}$, $\mathbb{P}_{A}$, $\mathbb{P}_{S}$ are given by $E_{c}(\mathbf{x}^{s})$, $E_{a}(\mathbf{x}^{s})$, and $E_{s}(\mathbf{x}^{s})$, respectively.
Sensitiveness loss. Since a sensitive factor is causally dependent on the sensitive attribute of data ($\mathbf{x}^{s},z^{s},y^{s}$), a simple classifier $h:\mathcal{A}\rightarrow\mathcal{Z}$ is learned, and further it is used to label the sensitive attribute in augmented data when learning $f$.
$\mathcal{L}_{sens}=CrossEntropy(z^{s},h(E_{a}(\mathbf{x}^{s})))$
Total loss. We jointly train the encoders and the decoder to optimize the transformation model $T$ with a weighted sum loss of the three terms.
$\min_{E_{c},E_{a},E_{s},G}\>\beta_{1}\mathcal{L}_{recon}^{data}+\beta_{2}\mathcal{L}_{recon}^{factor}+\beta_{3}\mathcal{L}_{sens}$ (3)
where $\beta_{1},\beta_{2},\beta_{3}>0$ are hyperparameters that control the importance of each loss term.
4.2 Fair Disentangled Domain Generalization
Furthermore, with a trained transformation model $T$, to learn the fairness-aware invariant classifier $f$ across domains, we make the following assumption.
@@ 4.2 Fair Disentangled Domain Generalization
We assume that inter-domain variation is characterized by covariate and dependence shifts. As a consequence, we assume that the conditional distribution $\mathbb{P}^{e}_{Y|XZ}$ is stable across domains, $\forall e\in\mathcal{E}$. Given a transformation model $T$, it holds that $\mathbb{P}^{e_{i}}_{Y|XZ}=\mathbb{P}^{e_{j}}_{Y|XZ},\>\forall e_{i},e_{j}\in\mathcal{E},i\neq j$, where $(X^{e_{j}},Z^{e_{j}})=T(\mathbf{X}^{e_{i}},Z^{e_{i}},e_{j})$.
In Assump. 2, the domain shift captured by $T$ would characterize the mapping from the marginal distributions $\mathbb{P}^{e_{i}}_{X}$ and $\rho(Y^{e_{i}},Z^{e_{i}})$ over $\mathcal{D}^{e_{i}}$ to the distribution $\mathbb{P}^{e_{j}}_{X}$ and $\rho(Y^{e_{j}},Z^{e_{j}})$ over $\mathcal{D}^{e_{j}}$ sampled from a different data domain $\mathbb{P}^{e_{j}}_{XZY}$, respectively. With this in mind and under Assump. 2, we introduce a new definition of fairness-aware invariance with respect to the variation captured by $T$ and satisfying the group fair constraint introduced in Defn. 1.
@@
Given a transformation model $T$, a fairness-aware classifier $f\in\mathcal{F}$ is domain invariant if it holds for all $e_{i},e_{j}\in\mathcal{E}$.
$f(\mathbf{x}^{e_{i}})=f(\mathbf{x}^{e_{j}}),\>\text{ and }\>\rho(f(X^{e_{i}}),Z^{e_{i}})=\rho(f(X^{e_{j}}),Z^{e_{j}})=0$ (4)
almost surely when $(\mathbf{x}^{e_{j}},z^{e_{j}})=T(\mathbf{x}^{e_{i}},z^{e_{i}},e_{j})$, $\mathbf{x}^{e_{i}}\sim\mathbb{P}^{e_{i}}_{X}$, $\mathbf{x}^{e_{j}}\sim\mathbb{P}^{e_{j}}_{X}$.
Defn. 3 is crafted to enforce invariance on the predictions generated by $f$ directly. We expect a prediction to remain consistent across various data realizations $T$ while considering group fairness.
Problem 2 (Fair Disentanglement Domain Generalization).
Under Defn. 3 and Assump. 2, if we restrict $\mathcal{F}$ of Prob. 1 to the set of invariant fairness-aware classifiers, the Prob. 1 is equivalent to the following problem
$P^{\star}\triangleq\min_{f\in\mathcal{F}}\>\>R(f)\triangleq\mathbb{E}_{\mathbb{P}^{s_{i}}_{XZY}}\ell(f(X^{s_{i}}),Y^{s_{i}})\quad\text{s.t.}\>\>f(X^{s_{i}})=f(X^{s_{j}}),\>\rho(f(X^{s_{i}}),Z^{s_{i}})=\rho(f(X^{s_{j}}),Z^{s_{j}})=0$ (5)
where $(X^{s_{j}},Z^{s_{j}})=T(X^{s_{i}},Z,s_{j})$, $\forall s_{i},s_{j}\in\mathcal{E}^{s}$, $i\neq j$.
Similar to [15], Prob. 2 is not a composite optimization problem. Moreover, acquiring domain labels is often expensive or even unattainable, primarily due to privacy concerns. Consequently, under the assumptions of disentanglement-based invariance and domain shift, Prob. 1 can be approximated to Prob. 2 by removing the max operator over $\mathcal{E}$.
In addition, Prob. 2 offers a new and theoretically-principled perspective on Prob. 1, when data varies from domain to domain with respect to $T$. To optimize Prob. 2 is challenging because (1) The strict equality constraints in Prob. 2 are difficult to enforce in practice; (2) Enforcing constraints on deep networks is known to be a challenging problem due to non-convexity. Simply transforming them to regularization cannot guarantee satisfaction for constrained problems; and (3) As we have incomplete access to all domains, it limits the ability to enforce fairness-aware $T$-invariance and further makes it hard to estimate $R(f)$.
@@
$P^{\star}(\gamma_{1},\gamma_{2})\triangleq\min_{f\in\mathcal{F}}R(f)\quad\text{s.t.}\quad\delta^{s_{i},s_{j}}(f)\leq\gamma_{1},\>\rho^{s_{i}}(f)\leq\frac{\gamma_{2}}{2},\>\text{and}\>\rho^{s_{j}}(f)\leq\frac{\gamma_{2}}{2}$ (6)
@@
$\delta^{s_{i},s_{j}}(f)$ $\triangleq\mathbb{E}_{\mathbb{P}^{s_{i}}_{XZ}}d\big{[}f(X^{s_{i}}),f(X^{s_{j}}=T(X^{s_{i}},Z^{s_{i}},s_{j}))\big{]},$ (7)
$\rho^{s_{i}}(f)$ $\triangleq\rho(f(X^{s_{i}}),Z^{s_{i}}),\quad\rho^{s_{j}}(f)\triangleq\rho(f(X^{s_{j}}),Z^{s_{j}})$ (8)
and $\forall s_{i},s_{j}\in\mathcal{E}^{s}$, $i\neq j$. Here, $\gamma_{1},\gamma_{2}>0$ are constants controlling the extent of relaxation and $d[\cdot]$ is a distance metric, e.g., KL-divergence. When $\gamma_{1}=\gamma_{2}=0$, Eqs. 5 and 6 are equivalent.
@@
$D^{\star}_{\xi,N,\mathcal{E}_{s}}(\gamma_{1},\gamma_{2})\triangleq$ $\max_{\lambda_{1}({s_{i},s_{j}}),\lambda_{2}({s_{i},s_{j}})}\min_{\boldsymbol{\theta}\in\Theta}\>\hat{R}(\boldsymbol{\theta})$
$+\frac{1}{|\mathcal{E}^{s}|}\sum_{{s_{i},s_{j}}\in\mathcal{E}_{s}}\Big{[}\lambda_{1}{(s_{i},s_{j})}\big{(}\hat{\delta}^{s_{i},s_{j}}(\boldsymbol{\theta})-\gamma_{1}\big{)}+\lambda_{2}{(s_{i},s_{j})}\big{(}\hat{\rho}^{s_{i}}(\boldsymbol{\theta})+\hat{\rho}^{s_{j}}(\boldsymbol{\theta})-\gamma_{2}\big{)}\Big{]}$ (9)
@@
where $\xi=\mathbb{E}_{\mathbb{P}_{X}}||f(\mathbf{x})-\hat{f}(\mathbf{x},\boldsymbol{\theta})||_{\infty}>0$ is a constant bounding the difference between $f$ and its parameterized counterpart $\hat{f}:\mathcal{X}\times\Theta\rightarrow\mathbb{R}$ defined in the Defn. 5.1 of [15]. $N$ is the number of samples drawn from $\mathbb{P}_{XZY}$ and it can be empirically replaced by $\sum_{s\in\mathcal{E}^{s}}|\mathcal{D}^{s}|$. $\lambda_{1}{(s_{i},s_{j})},\lambda_{2}{(s_{i},s_{j})}>0$ are dual variables. $\hat{R}(\boldsymbol{\theta})$, $\hat{\delta}^{s_{i},s_{j}}(\boldsymbol{\theta})$, $\hat{\rho}^{s_{i}}(\boldsymbol{\theta})$ and $\hat{\rho}^{s_{j}}(\boldsymbol{\theta})$ are the empirical counterparts of $R(\hat{f}(\cdot,\boldsymbol{\theta}))$, $\delta^{s_{i},s_{j}}(\hat{f}(\cdot,\boldsymbol{\theta}))$, $\rho^{s_{i}}(\hat{f}(\cdot,\boldsymbol{\theta}))$ and $\rho^{s_{j}}(\hat{f}(\cdot,\boldsymbol{\theta}))$, respectively.
Algorithm 1 FEDORA: Fair Disentangled Domain Generalization.
Require: Encoders $E=\{E_{c},E_{a},E_{s}\}$, decoder $G$ and sensitive classifier $h$.
@@
3:         $\mathcal{L}_{cls}(\boldsymbol{\theta})=\frac{1}{m}\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta}))$
4:         Initialize $\mathcal{L}_{inv}(\boldsymbol{\theta})=0$ and $\mathcal{B}^{\prime}=[\>]$
@@
Generate $(\mathbf{x}_{j},z_{j},y_{j})=T(\mathbf{x}_{i},z_{i},y_{i})$ and add to $\mathcal{B}^{\prime}$
@@
$\mathcal{L}_{inv}(\boldsymbol{\theta})+=\frac{1}{m}d[\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta}),\hat{f}(\mathbf{x}_{j},\boldsymbol{\theta})]$